The following transactions pertain to Accounting Solutions Inc. Assume the transactions for the purchase of the computer and any capital improvements occur on January 1 each year. Year 1 Acquired $70,000 cash from the issue of common stock. Purchased a computer system for $22,800. It has an estimated useful life of five years and a $2,520 salvage value. Paid $1,900 sales tax on the computer system. Collected $34,590 in fees from clients. Paid $1,050 in fees for routine maintenance to service the computers. Recorded double-declining-balance depreciation on the computer system for Year 1. Year 2 Paid $960 for repairs to the computer system. Bought off-site backup services to maintain the computer system, $1,160. Collected $37,590 in fees from clients. Paid $910 in fees to service the computers. Recorded double-declining-balance depreciation for Year 2. Year 3 Paid $2,700 to upgrade the computer system, which extended the total life of the system to six years. The salvage value did not change. Paid $830 in fees to service the computers. Collected $37,890 in fees from clients. Recorded double-declining-balance depreciation for Year 3. Required a. Record the above transactions in a horizontal statements model. b-1. Prepare income statements for Year 1, Year 2, and Year 3. b-2. Prepare statements of changes in stockholders' equity for Year 1, Year 2, and Year 3. b-3. Prepare balance sheets for Year 1, Year 2, and Year 3. b-4. Prepare statements of cash flows for Year 1, Year 2, and Year 3.
Answer #1
a. Accounting solutions Inc.
Horizontal statements model-Year 1
Event Year 1 Balance sheet Income statement Statement of cash flows
Assets = Stockholder's equity Revenue - Expense = Net income
Cash + Equipment - Accumulated depreciation = Common stock + Retained earnings
1 70000 70000 70000 FA
2 -22800 22800 -22800 IA
3 -1900 1900 -1900 IA
4 34590 34590 34590 0 34590 34590 OA
5 -1050 -1050 0 1050 -1050 -1050 OA
6 9880 -9880 0 9880 -9880 0 NA
Bal. 78840 24700 9880 70000 23660 34590 10930 23660
6
Depreciation rate=2*(1/Useful life)=2*(1/5)=2*0.20=0.40=40%
Depreciation expense=Book value at the end of year 1*Depreciation rate
Book value at the end of year 1=22800+1900=$ 24700
Depreciation expense=24700*40%=$ 9880
Accounting solutions Inc.
Horizontal statements model-Year 2
Event Year 2 Balance sheet Income statement Statement of cash flows
Assets = Stockholder's equity Revenue - Expense = Net income
Cash + Equipment - Accumulated depreciation = Common stock + Retained earnings
Bal. 78840 24700 9880 70000 23660
1 -960 -960 0 960 -960 -960 OA
2 -1160 -1160 0 1160 -1160 -1160 OA
3 37590 37590 37590 0 37590 37590 OA
4 -910 -910 0 910 -910 -910 OA
5 5928 -5928 0 5928 -5928 0 NA
Bal. 113400 24700 15808 70000 52292 37590 8958 28632
5
Depreciation expense=Book value at the end of year 2*Depreciation rate
Book value at the end of year 2=Cost-Accumulated depreciation=24700-9880=$ 14820
Depreciation expense=14820*40%=$ 5928

Event Year 3 Bal. Statement of cash flows Cash Accounting solutions Inc. Horizontal statements model-Year 3 Balance sheet Inca 34590 37590 37890 Service revenue Expenses Depreciation expense Maintenance expense Service fee expense 9880 4637 1050 5928For the year ended year 1 Year 1 Year 2 Year 3 Assets 78840 113400 147760 24700 24700 15808 9880 24700 17745 154715 93660 122Cash flows from investing activities Cash outflow for equipment b -24700 0 -2700 Cash flows from financing activities Cash in